3x-y=11
5y-7x=1
I have to find the x and y. Thanks for your help.

Found 2 solutions by stanbon, TigerAlgebra:
Answer by stanbon(75887)

3x-y=11
5y-7x=1
---------
Modify for elimination:
-5y + 15x = 55
5y - 7x = 1
------------------
Add and solve for "x":
8x = 56
x = 7
---------
Solve for "y":
3x-y = 11
21 - y = 11
y = 10
---------------
Cheers,
Stan H.
----------------

Answer by TigerAlgebra(40)